# From Pen to Prompt: How Creative Writers Integrate AI into their Writing Practice

### v2

@@ Abstract.
Creative Writing, Large Language Model, Human-AI Collaboration
$c$s: Human-centered computing Collaborative and social computing systems and tools
@@ 1. Introduction
But the unique properties of creative writing—as opposed to academic, business, journalistic, or marketing writing, where LLMs have proven to help with productivity (Jin et al., 2024; Zhang et al., 2021; Jovic and Mnasri, 2024; Zhou et al., 2024)—raise additional concerns about how AI usage can impact aspects of one’s writing craft, like authenticity, ownership, and the creative process (Hwang et al., 2024; Gero et al., 2023). For some, AI feels fundamentally incompatible with the creative act, reducing a multitude of creative choices to a few prompts (Chiang, 2024). Others view AI as potentially another tool in a writer’s creative arsenal, akin to digital word processors and editing software (Biermann et al., 2022). This tension is evident as literary magazines and creative writing platforms have struggled to decide on a policy for AI-assisted work amidst strong disagreements in the creative writing community (Weatherbed, 2024).
Previous work has explored writers’ hypothetical responses to AI and studied the use of specific tools through lab experiments (Calderwood et al., 2020; Chakrabarty et al., 2024; Singh et al., 2023; Ippolito et al., 2022). However, few studies have examined the long-term qualitative experience of creative writers who have actively chosen to integrate AI into their workflows and its effects on both their craft and career. These writers suggest that there may be ways to preserve essential human elements of creative writing while embracing new technological possibilities.
@@
Our participants represent a distinct group: creative writers who have chosen to use AI despite potential concerns. In this moment where the role of AI in creative work is still evolving, these writers are uniquely positioned to reflect on their experiences before and after writing with AI. Having used AI long enough to develop their own practices, they shed light on sustainable approaches to AI integration in creative practice. Through interviews and direct observation of their writing processes, we examine their strategies used to navigate tensions around AI and the emergent practices that arise over continued use.
We find that creative writers actively negotiate between AI’s benefits with their deeply held writerly values, setting boundaries around AI use to preserve love of the craft. Rather than viewing AI simply as tool or collaborator, writers develop fluid relationships with AI that shift with their goals, influencing and reinforcing AI use. These relationships are supported by sophisticated workflows involving many iterations and decision points, maintaining creative control through careful decision-making at every stage. We characterize this dynamic interplay of values, relationships, and strategies to reveal how different approaches to AI emerge across individuals and genres. Our findings provide insight for writing communities, tool designers, and future research. As the landscape evolves, understanding how writers maintain their craft and values while embracing technological support will be crucial for developing AI and practices that support instead of undermine human creativity.
@@ 2. Related Work
To understand why and how creative writers are integrating AI into their practice, we look at the current issues of AI in creative writing communities, along with previous work in computational support for creative writing and studies of their writer-AI interactions.
@@ 2.1. Effects of AI on creative writing communities
While AI benefits individual writers by helping them be more productive and creative(Doshi and Hauser, 2024; Noy and Zhang, 2023), studies also suggest AI can lead to creative homogenization (Anderson et al., 2024; Li et al., 2024a; Doshi and Hauser, 2024) . Self-publishing platforms have seen a surge in AI-assisted work (Acovino, 2023; Sato, 2023), simultaneously raising concerns about market saturation and spam. Meanwhile, tools like “Have I Been Trained” (AI, 2025) showing how writers’ works are being trained on without consent have fueled demands for better frameworks around data rights, compensation, and legal clarity around copyright. Creative institutions are still adapting—some magazines temporarily halted submissions after being inundated with AI generated content largely from outside the writing community (Acovino, 2023), and many venues have introduced explicit policies banning or requiring disclosure of AI use. The decision of NaNoWriMo (a popular month-long novel writing challenge) to allow AI use faced intense backlash, highlighting divisions within creative writing communities (Weatherbed, 2024).
@@ 2.2. Computational support for creative writing
Computational tools have supported creative writing for decades, from rule-based systems(Howe, 2009; Compton and Mateas, 2015b) to chatbots. Over time, computational support evolved to address almost every stage of the writing process from ideation (Shaer et al., 2024; Han et al., 2018; Wan et al., 2024) to rewriting and feedback (Chakrabarty et al., 2024; Zhou and Sterman, 2024) across creative genres ranging from storytelling (Kreminski et al., 2020) to poetry (Gero and Chilton, 2019b) to theater scripts (Mirowski et al., 2023). Modern AI writing systems offer integration at multiple stages of the writing process. For example, commercial AI platforms made for creative writers, such as Sudowrite (Sud, [n. d.]) and Novelcrafter (nov, [n. d.]) generate inline text continuations that maintain stylistic and narrative consistency alongside more focused features such as structured story outlining and worldbuilding assistance. Chat-based platforms like ChatGPT(cha, [n. d.]) let writers explicitly prompt for creative writing support. Other tools provide support for focused tasks, such as generating customized thesauruses (Gero and Chilton, 2019a), suggesting connections for new metaphors (Gero and Chilton, 2019b), and building characters through chat (Qin et al., 2024) or simulation (Kreminski et al., 2020).
@@ 2.3. Understanding writer interactions with AI
A growing body of work has characterized writer perceptions and usage of AI, with writers finding AI helpful in the planning, translating, and reviewing stages of the writing process (Calderwood et al., 2020; Chakrabarty et al., 2024; Singh et al., 2023; Benharrak et al., 2024); however, much of this work uses controlled lab studies with custom AI interfaces. Differences between writers emerged over longer settings—when asked to use a custom AI interface over an 8-week period, researchers observed some choosing to incorporate AI text directly while others used it mainly for ideas, alongside common uses such as continuation elaborating, replacing, rewriting (Ippolito et al., 2022). Studies also observed differences between types of writers. Novices perceive greater benefits from using AI for ideation while experienced writers struggle with aligning AI-generated content to their vision (Li et al., 2024b; Ippolito et al., 2022). Hobbyists emphasize personal expression and ownership, while professionals prioritize productivity (Biermann et al., 2022). Indeed, in broader creative domains, researchers have seen a shift in creative professionals’ work towards high-level orchestration much like that of a project managers’ workflow (Palani and Ramos, 2024).
@@
Relatively less research has examined interaction paradigms between writers and AI. Work characterizing writing support tools found that they can be categorized anywhere from cognitive aids to collaborators (Gero and Chilton, 2019b; Compton and Mateas, 2015a; Yeh et al., 2024; Yang et al., 2022; Ghajargar et al., 2022). To understand the interplay between writer and AI, researchers have built visualizations and analyses of log data capturing writer interactions with AI from crowdsourced participants (Lee et al., 2022; Shibani et al., 2023). As creative writers interact with AI tools, common challenges include preserving narrative coherence (Ippolito et al., 2022; Li et al., 2024b), maintaining stylistic voice (Draxler et al., 2023; Ippolito et al., 2022), diminishing senses of ownership and satisfaction (Hwang et al., 2024; Kobiella et al., 2024), and worries about plagiarism (Göldi et al., 2024). Moreover, depending on its use, AI assistance does not necessarily improve text quality (Göldi et al., 2024; Clark et al., 2018). Finally, social dynamics impact when writers might prefer AI support or human peers, with factors including writers’ desires for support, perceptions of support actors, and values about writing (Gero et al., 2023).
@@
This work builds upon these foundations by examining writers who have chosen to incorporate AI into their creative practice and have had months to years to develop their workflows. Previous approaches often fragment the creative process by focusing on specific stages or tools, with participants who are either new to AI for creative writing or the tools they are asked to use. Research has shown that AI tools become more useful over time as users familiarize themselves, highlighting the importance of sustained engagement with tools rather than experimental use (Long et al., 2024). Rather than studying specific tools or writing tasks in isolation, we observe writers’ complete workflows across multiple commercial AI writing platforms from different genres. By combining interviews with direct observation of writers’ practices, we capture the concrete decisions and emergent patterns that develop in practice as writers establish sustainable approaches to working with AI.
@@ 3. Methods
Over the course of November 2023 to July 2024, the first author conducted 90-minute semi-structured interviews virtually with the writers, spending 30 minutes of the interview time directly observing their writing process.
@@ 3.1. Recruitment and Participants
The final participants included 18 creative writers in different genres, with fiction most represented (Table 1). Writers’ experience with AI span months to years, and they use a variety of AI tools spanning chatbots to more integrated inline-generation platforms with many focused features. Writers’ use of AI is detailed in Appendix B. Nine had no formal writing education, 3 had masters/grad degrees in writing, and 6 took writing classes in college or online. Twelve participants resided in the USA, 5 in Europe, 1 in Canada, and 1 undisclosed. Ten participants identified as female, 6 as male, 1 as non-binary, and 1 undisclosed.
@@ 3.2. Interview Procedure
Interviews included an observed live-writing session lasting 30–45 minutes, where writers shared their screen and worked on an active writing project using their preferred AI platforms. Participants were instructed to think aloud while writing as they normally would. The researcher asked contextual questions about specific actions such as “what made you accept this suggestion?” Writing while being observed can be a potentially uncomfortable position. To minimize observer effects, participants were given alternatives to live observation: 1) prerecording their writing process before the interview, or 2) keeping a diary of their writing process before the interview. One participant prerecorded themselves writing, then participated in the interview where the writer asked questions about their recording in place of the observation session. The final interview guideline is detailed in Appendix A. This study was reviewed by our institution’s IRB and deemed exempt.
@@ 3.3. Data Analysis
We analyzed interview and writing session data following grounded theory methods (Muller and Kogan, 2012). During data collection, the first author took detailed notes, and the interview recordings were automatically transcribed. Initial codes were developed by the first and third author, then updated with the second author as new interviews were conducted. We additionally coded non-verbal interactions (e.g., prompting method) from the screen recordings that transcripts couldn’t capture. The authors discussed emerging themes after every few interviews and updated the interview guideline accordingly. The data was open-coded by the authors and clustered into themes, with the first author leading the process.
@@ 4. Findings
We identified three key elements that shape how writers engage with AI: their personal values about writing, their perceived relationships with AI, and their concrete AI integration strategies. Rather than operate independently, these elements form a dynamic system where each element influences and reinforces the others. For example, writers who strongly value ownership will often relate to AI as an assistant or tool, in order to maintain explicit control. These relationships in turn shape what integration strategies writers develop: writers viewing AI as an assistant typically develop more bounded, task-specific workflows compared to those who see AI as a collaborator. However, influence also flows in the other direction, where writers’ experiences with using AI can shift how they view their relationship over time.
@@ 4.1. When and Why Do Creative Writers Use (and Not Use) AI?
We found that values play a major role in when and why writers choose to use AI, setting usage boundaries to preserve what matters most in their creative practice. Thus, we begin by briefly summarizing the values writers expressed in our interviews. While prior work has explored taxonomies of writers’ values that overlap with our findings (Ippolito et al., 2022; Gero et al., 2023; Biermann et al., 2022), the goal of our inquiry is to understand how such factors shape writers’ engagement with AI. To ground our further discussion of findings, the core writers’ values that emerged include:
@@
A fuller description of these values and example quotes can be found in Appendix C.
4.1.1. What motivates writers to use AI?
@@
AI redefines relationships with creative blocks. The most common benefit was that AI transforms writer’s block from an intimidating barrier into a manageable challenge. AI provided starting points as ideas or text, reducing the intimidation and cognitive load of figuring out what comes next, “where with the AI, you don’t have the mental game tripping you up anymore” (P2). Writers don’t always expect high-quality responses, but maintained confidence in their ability to shape any AI output into something useful. P13 adds: “I have a very low standard/expectation for AI’s generative responses, so I usually just take what it gives me, and I usually can run with that.” AI’s ability to offer concrete ways to move forward acts as a “safety net”, helping writers maintain creative momentum and finish their writing.
@@
AI can be a productivity boost and accessibility tool. AI can accelerate parts of the writing process, such as ideation and editing, where AI’s constant availability shortens feedback loops. The meaning of “productivity” varied among participants. For writers relying on writing as a source of income, productivity is tied to maintaining competitive publishing speeds in an increasingly demanding market. For hobbyists, AI allows writers to complete projects otherwise constrained by time or energy. For writers managing health challenges, productivity shifted to enabling participation. P7, who struggles with carpal tunnel, uses AI to generate prose during first drafts: “I do have flare ups… I was even considering hiring a co-author or ghostwriter… To have the AI do that really helped, really made me think that I can actually write again…” P14, who writes for a living, mentioned AI as leveling the publishing playing field: “As someone who has had physical disabilities and the inability to work 40 hours a week for most of my adult life, suddenly I am competitive…
@@
Writers can do more of what they love while developing their craft. Writers use AI strategically to reduce friction in challenging areas while maintaining focus on the parts of writing they find most fulfilling, creating a more satisfying creative practice. Each writer identified different parts of the writing process they enjoy or struggle with (such as plotting, writing dialogue, character descriptions). Writers also noted that working with AI for challenging aspects became a learning opportunity, helping them develop their skills through example: “this is a good learning experience for me too, because I’m not a perfect songwriter, I’m not a perfect poet. I don’t think anybody is” (P13).
@@
AI enables new forms of creative experimentation. Writers deeply enjoy being creative, while still conceding “everything’s already been done before” (P12). For many, creativity is instead more about the personal process of experimentation and shaping ideas in unique ways. Writers are looking for new ways of using AI beyond a basic writing helper and enjoy creatively experimenting in a low-stakes environment. For example, P3 noted wanting “smaller, less fluent models producing random but cooler, wackier text… I would love to experiment with that.” As another example, AI can help writers venture beyond their usual writing styles and genres: “within a lifetime, you can write more work, write more types of work, or you can be exposed to a lot of different genres really easily” (P17).
4.1.2. When are writers cautious about using AI?

### v4

@@ Abstract.
Creative Writing, Large Language Model, Human-AI Collaboration, AI writing support, Personal Values
journalyear: 2025copyright: ccconference: Creativity and Cognition; June 23–25, 2025; Virtual, United Kingdombooktitle: Creativity and Cognition (C&C ’25), June 23–25, 2025, Virtual, United Kingdomdoi: 10.1145/3698061.3726910isbn: 979-8-4007-1289-0/2025/06ccs: Human-centered computing Empirical studies in HCIccs: Applied computing Arts and humanities
@@ 1. Introduction
But the unique properties of creative writing—as opposed to academic, business, journalistic, or marketing writing, where LLMs have proven to help with productivity (Jin et al., 2024; Zhang et al., 2021; Jovic and Mnasri, 2024; Zhou et al., 2024)—raise additional concerns about how AI usage can impact aspects of one’s writing craft, like authenticity, ownership, and the creative process (Hwang et al., 2024; Gero et al., 2023). For some, AI feels fundamentally incompatible with the creative act, reducing a multitude of creative choices to a few prompts (Chiang, 2024). Others view AI as potentially another tool in a writer’s creative arsenal, akin to digital word processors and editing software (Biermann et al., 2022). This tension is evident as literary magazines and creative writing platforms have struggled to decide on policies for AI-assisted work amidst strong disagreements in the creative writing community (Weatherbed, 2024).
Previous work has explored writers’ hypothetical responses to AI and studied the use of specific tools through lab experiments (Calderwood et al., 2020; Chakrabarty et al., 2024a; Singh et al., 2023; Ippolito et al., 2022). Many of these studies primarily focus on isolated writing tasks or specific AI tools in controlled settings, rather than observing writers’ complete practices across multiple platforms. Additionally, prior studies often involve participants who are new to AI writing tools, capturing initial reactions rather than evolved practices. This prior research establishes important foundations regarding writers’ initial concerns and challenges with AI tools—however, few studies have examined the long-term qualitative experience of creative writers who have actively chosen to integrate AI into their workflows and its effects on both their craft and career. This experience becomes especially important as research has shown that AI tools become more useful over time as users familiarize themselves (Long et al., 2024). These writers suggest that there may be ways to preserve essential human elements of creative writing while embracing new technological possibilities.
@@
Our participants represent a distinct group: creative writers who have chosen to use AI despite potential concerns. In this moment where the role of AI in creative work is still evolving, these writers are uniquely positioned to reflect on their experiences before and after writing with AI. Having used AI long enough to develop their own practices, they shed light on sustainable approaches to AI integration in creative practice. Through interviews and direct observation of their writing processes, we examine their strategies used to navigate tensions around AI and the emergent practices that only arise over continued use.
We find that creative writers actively negotiate between AI’s benefits and their deeply held writerly values, setting boundaries around AI use to preserve love of the craft. Rather than viewing AI simply as tool or collaborator, writers develop fluid relationships with AI that shift with their goals, influencing and reinforcing AI use. These relationships are supported by sophisticated workflows involving many iterations and decision points, maintaining creative control through careful decision-making at every stage. We characterize this dynamic interplay of values, relationships, and strategies to reveal how different approaches to AI emerge across individuals and genres. Our findings provide insight for writing communities, tool designers, and future research. As the landscape evolves, understanding how writers maintain their craft and values while embracing technological support will be crucial for developing AI and practices that support instead of undermine human creativity.
@@ 2. Related Work
To understand why and how creative writers are integrating AI into their practice, we look to current issues with AI in creative writing communities, along with previous work in computational support for creative writing and studies of writer-AI interactions.
@@ 2.1. Effects of AI on creative writing communities
While AI benefits individual writers by helping them be more productive and creative (Doshi and Hauser, 2024; Noy and Zhang, 2023), studies also suggest AI can lead to creative homogenization (Anderson et al., 2024; Li et al., 2024a; Doshi and Hauser, 2024). Self-publishing platforms have seen a surge in AI-assisted work (Acovino, 2023; Sato, 2023), simultaneously raising concerns about market saturation and spam. Meanwhile, tools like “Have I Been Trained” (Spawning AI, 2025) showing how writers’ works are being trained on without consent have fueled demands for better frameworks around data rights, compensation, and legal clarity around copyright. Creative institutions are still adapting—some magazines temporarily halted submissions after being inundated with AI generated content largely from outside the writing community (Acovino, 2023), and many venues have introduced explicit policies banning or requiring disclosure of AI use. The decision of NaNoWriMo (a popular month-long novel writing challenge) to allow AI use faced intense backlash, highlighting divisions within creative writing communities (Weatherbed, 2024).
@@ 2.2. Computational support for creative writing
Computational tools have supported creative writing for decades, from rule-based systems (Howe, 2009; Compton and Mateas, 2015b) to chatbots. Over time, computational support evolved to address almost every stage of the writing process from ideation (Shaer et al., 2024; Han et al., 2018; Wan et al., 2024) to rewriting and feedback (Chakrabarty et al., 2024b; Zhou and Sterman, 2024) across creative genres ranging from storytelling (Kreminski et al., 2020) to poetry (Gero and Chilton, 2019b) to theater scripts (Mirowski et al., 2023). Modern AI writing systems offer integration at multiple stages of the writing process. For example, commercial AI platforms made for creative writers, such as Sudowrite (Sud, [n. d.]) and Novelcrafter (nov, [n. d.]) generate inline text continuations that maintain stylistic and narrative consistency alongside more focused features such as structured story outlining and worldbuilding assistance. Chat-based platforms like ChatGPT (cha, [n. d.]) let writers explicitly prompt for creative writing support. Other tools provide support for focused tasks, such as generating customized thesauruses (Gero and Chilton, 2019a), suggesting connections for new metaphors (Gero and Chilton, 2019b), and building characters through chat (Qin et al., 2024) or simulation (Kreminski et al., 2020).
@@ 2.3. Understanding writer interactions with AI
A growing body of work has characterized writer perceptions and usage of AI, with writers finding AI helpful in the planning, translating, and reviewing stages of the writing process (Calderwood et al., 2020; Chakrabarty et al., 2024a; Singh et al., 2023; Benharrak et al., 2024); however, much of this work uses controlled lab studies with custom AI interfaces. Differences between writers emerged over longer settings—when asked to use a custom AI interface over an 8-week period, researchers observed some choosing to incorporate AI text directly while others chose to use AI primarily for ideation and rewriting (Ippolito et al., 2022). Studies also observed differences between types of writers. Novices perceive greater benefits from using AI for ideation while experienced writers struggle with aligning AI-generated content to their vision (Li et al., 2024b; Ippolito et al., 2022). Hobbyists emphasize personal expression and ownership, while professionals prioritize productivity (Biermann et al., 2022). Indeed, in broader creative domains, researchers have seen a shift in creative professionals’ work towards high-level orchestration much like that of a project managers’ workflow (Palani and Ramos, 2024).
@@
Relatively less research has examined interaction paradigms between writers and AI. Work characterizing writing support tools found that they can be categorized anywhere from cognitive aids to collaborators (Gero and Chilton, 2019b; Compton and Mateas, 2015a; Yeh et al., 2024; Yang et al., 2022; Ghajargar et al., 2022). To understand the interplay between writer and AI, researchers have built visualizations and analyses of log data capturing writer interactions with AI from crowdsourced participants (Lee et al., 2022; Shibani et al., 2023). As creative writers interact with AI tools, common challenges include preserving narrative coherence (Ippolito et al., 2022; Li et al., 2024b), maintaining stylistic voice (Draxler et al., 2023; Ippolito et al., 2022), diminishing senses of ownership and satisfaction (Hwang et al., 2024; Kobiella et al., 2024), and worries about plagiarism (Göldi et al., 2024). Moreover, depending on its use, AI assistance does not necessarily improve text quality (Göldi et al., 2024; Clark et al., 2018). Finally, social dynamics impact when writers might prefer AI support to a human peer, with factors including writers’ desires for support, perceptions of human or AI support, and values about writing (Gero et al., 2023).
@@
This work builds upon these foundations by examining writers who have chosen to incorporate AI into their creative practice and have had months to years to develop their workflows. Previous approaches often fragment the creative process by focusing on specific stages or tools, with participants who are either new to AI for creative writing or the tools they are asked to use. Research has shown that AI tools become more useful over time as users learn and familiarize themselves with the system, move past the novelty effect, and begin customizing and appropriating the technology (Long et al., 2024; Koch et al., 2018), highlighting the importance of sustained engagement with tools rather than exploratory use. Rather than studying specific tools or writing tasks in isolation, we observe writers’ complete workflows across multiple commercial AI writing platforms from different genres. By combining interviews with direct observation of writers’ practices, we capture the concrete decisions and emergent patterns that only develop in practice as writers establish sustainable approaches to working with AI past the familiarization phase.
@@ 3. Methods
Over the course of November 2023 to July 2024, the first author conducted 90-minute semi-structured interviews virtually with the writers, spending around 30 minutes of the interview time directly observing their writing process.
@@ 3.1. Recruitment and Participants
The final participants included 18 creative writers in different genres, with fiction most represented (Table 1). Writers’ experience with AI span months to years, and they use a variety of AI tools from chatbots to more integrated inline-generation platforms with many focused features. Writers’ reported use of AI is detailed in Appendix B. Nine had no formal writing education, 3 had masters/grad degrees in writing, and 6 took writing classes in college or online. Twelve participants resided in the USA, 5 in Europe, 1 in Canada, and 1 undisclosed. Ten participants identified as female, 6 as male, 1 as non-binary, and 1 undisclosed.
@@ 3.2. Interview Procedure
Interviews included an observed live-writing session lasting approximately 30 minutes (range: 29-45 minutes, mean: 34.5 minutes), where writers shared their screen and worked on a current writing project (novel chapters, poems, essays) using their preferred AI platforms (see Appendix E for details on participants’ writing projects and setups). Participants were instructed to write naturally as they would in their normal practice, without any constraints on their process or AI usage and encouraged to think aloud about their decision-making to the degree they felt comfortable. The researcher asked contextual questions about specific actions such as “what made you accept this suggestion?” but otherwise minimized intervention to preserve the authenticity of their process. Writing while being observed can be a potentially uncomfortable position. To minimize observer effects, participants were given alternatives to live observation: 1) prerecording their writing process before the interview, or 2) keeping a diary of their writing process before the interview. One participant chose option 1 and prerecorded themselves writing, then participated in the interview where the writer asked questions about their recording in place of the observation session. The final interview guideline is detailed in Appendix A. This study was reviewed by our institution’s IRB and deemed exempt.
@@ 3.3. Data Analysis
We analyzed interview and writing session data following grounded theory methods (Muller and Kogan, 2012). After data collection, the interview recordings were automatically transcribed and aligned with the videos, then manually corrected for errors. Initial codes were developed by the first and third author, then updated with the second author as new interviews were conducted. We additionally coded non-verbal interactions (e.g., prompting method) from the screen recordings that transcripts couldn’t capture. The authors discussed emerging themes after every few interviews and updated the interview guideline accordingly. The data was open-coded by the authors and clustered into themes, with the first author leading the process.
@@ 4. Findings
We identified three key elements that shape how writers engage with AI: their personal values about writing, their perceived relationships with AI, and their concrete AI integration strategies. Rather than operate independently, these elements form a dynamic system where each element influences and reinforces the others. For example, writers who strongly value ownership will often relate to AI as an assistant or tool, in order to maintain explicit control. These relationships in turn shape what integration strategies writers develop: writers viewing AI as an assistant typically develop more bounded, task-specific workflows compared to those who see AI as a collaborator.
@@ 4.1. When and Why Do Creative Writers Use (and Not Use) AI?
We found that values play a major role in when and why writers choose to use AI, setting usage boundaries to preserve what matters most in their creative practice. To ground our findings, we begin by briefly summarizing the core writers’ values that emerged:
@@
A fuller description of these values and example quotes can be found in Appendix D. While prior work has explored taxonomies of writers’ values and writer challenges that overlap with our findings (Ippolito et al., 2022; Gero et al., 2023; Biermann et al., 2022), the goal of our inquiry is to understand how such factors shape writers’ engagement with AI. We present our results in relation to this earlier work, highlighting how our sample of experienced AI users extends existing insights. The values of intention, authenticity, and creativity from Gero et al. in the context of social dynamics characterize when writers would choose to seek help from AI rather than a peer (Gero et al., 2023). Biermann et al. distinguishes the emotional values of fulfillment, ownership, and integrity, in contrast with the value of productivity as competing motivations in writers’ willingness to use AI. Building on this prior work, our study examines not only when and why writers choose to use AI, but also how their values guide the way they use it. Prior challenges associated with AI such as maintaining writer voice, avoiding tropes, and understanding story structures (Ippolito et al., 2022) are somewhat mitigated in our study due to the advancement in model capabilities, the participants’ experience with AI platforms, and the use of commercial platforms specifically designed for creative writing.
4.1.1. What motivates writers to use AI (and keep using it)?
@@
AI redefines relationships with creative blocks. The most common benefit was that AI transforms writer’s block from an intimidating barrier into a manageable challenge. AI provided starting points as ideas or text, reducing the intimidation and cognitive load of figuring out what comes next, “where with the AI, you don’t have the mental game tripping you up anymore” (P2). Writers don’t always expect high-quality responses, but maintained confidence in their ability to shape any AI output into something useful. P13 adds: “I have a very low standard/expectation for AI’s generative responses, so I usually just take what it gives me, and I usually can run with that.” AI’s ability to offer concrete ways to move forward acts as a “safety net,” helping writers maintain creative momentum and finish their writing.
@@
AI can be a productivity boost and accessibility tool. AI can accelerate parts of the writing process, such as ideation and editing, where AI’s constant availability shortens feedback loops. Biermann et al. (2022) distinguish between hobbyist writers who resist AI to preserve emotional values and see AI as a tool, and professionals who prioritize productivity and see AI as a writing collaborator (Biermann et al., 2022). We find that the meaning of “productivity” varied among participants. For writers relying on writing as a source of income (P2, P10, P12, P14), productivity is tied to maintaining competitive publishing speeds in an increasingly demanding market. For hobbyists, AI allows writers to complete projects otherwise constrained by time or energy for which some welcomed AI help in translating their ideas to prose and still found ways to get fulfillment from expressing their ideas. For writers managing health challenges, productivity shifted to enabling participation. P7, who struggles with carpal tunnel, uses AI to generate prose during first drafts: “I do have flare ups… I was even considering hiring a co-author or ghostwriter… To have the AI do that really helped, really made me think that I can actually write again…” P14, who writes for a living, mentioned AI as leveling the publishing playing field: “As someone who has had physical disabilities and the inability to work 40 hours a week for most of my adult life, suddenly I am competitive…
@@
Writers can do more of what they love while developing their craft. Writers use AI strategically to reduce friction in challenging areas while maintaining focus on the parts of writing they find most fulfilling, creating a more satisfying creative practice. Each writer identified different parts of the writing process they enjoy or struggle with (such as plotting, writing dialogue, character descriptions). Writers, especially hobbyists, also noted that working with AI for challenging aspects became a learning opportunity, helping them develop their skills through example: “this is a good learning experience for me too, because I’m not a perfect songwriter, I’m not a perfect poet. I don’t think anybody is” (P13). While prior work suggests that writers are more open to AI collaboration in areas where they lack confidence, our findings complicate that view (Biermann et al., 2022). Writers in our study used AI strategically to reduce friction in parts of the process they found less enjoyable and not necessarily those they were less skilled in, though the two are often related. In fact, several experienced writers (especially in genre fiction and poetry) used AI in areas where they felt confident but felt to be tedious. Conversely, some hobbyist essayists avoided AI even in areas they struggled with, due to concerns about voice or authenticity.
@@
AI enables new forms of creative experimentation. Writers deeply enjoy being creative, while still conceding “everything’s already been done before” (P12). For many, creativity is instead more about the personal process of experimentation and shaping ideas in unique ways. Writers are looking for new ways of using AI beyond a basic writing helper and enjoy creatively experimenting in a low-stakes environment, especially if they do not wish to use AI for prose generation directly. For example, P3, who enjoyed creating sentences and did not want to use any full sentences generated by AI noted wanting “smaller, less fluent models producing random but cooler, wackier text… I would love to experiment with that.” As another example, AI can help writers venture beyond their usual writing styles and genres: “within a lifetime, you can write more work, write more types of work, or you can be exposed to a lot of different genres really easily” (P17).
4.1.2. When are writers cautious about using AI in their writing?